# Imputation-free Learning of Tabular Data with Missing Values using Incremental Feature Partitions in Transformer

### v4

@@ Appendix B Data set-specific AUC classification scores
Dataset ID % Median-GBT Median-FTT MICE-GBT MICE-FTT GAIN-GBT GAIN-FTT Diffputer-GBT Diffputer-FTT MIA-Xgboost MIA-LightGBM AM-FTT NIIL (K=d/2)
1063 10 0.892 (0.05) 0.623 (0.11) 0.863 (0.04) 0.668 (0.14) 0.496 (0.03) 0.509 (0.02) 0.468 (0.06) 0.525 (0.11) 0.904 (0.04) 0.912 (0.04) 0.783 (0.07) 0.871 (0.05)
50 0.553 (0.06) 0.536 (0.05) 0.506 (0.04) 0.499 (0.05) 0.505 (0.02) 0.512 (0.03) 0.507 (0.08) 0.486 (0.05) 0.540 (0.07) 0.542 (0.08) 0.492 (0.04) 0.557 (0.03)
463 10 0.627 (0.09) 0.675 (0.06) 0.596 (0.11) 0.485 (0.12) 0.672 (0.10) 0.627 (0.22) 0.532 (0.08) 0.541 (0.21) 0.665 (0.08) 0.595 (0.10) 0.406 (0.09) 0.746 (0.06)
50 0.652 (0.13) 0.595 (0.12) 0.618 (0.14) 0.439 (0.12) 0.557 (0.09) 0.595 (0.15) 0.663 (0.09) 0.563 (0.11) 0.623 (0.10) 0.543 (0.11) 0.602 (0.21) 0.657 (0.10)
13 10 0.704 (0.08) 0.657 (0.08) 0.704 (0.08) 0.645 (0.06) 0.701 (0.09) 0.527 (0.13) 0.666 (0.08) 0.694 (0.06) 0.698 (0.08) 0.655 (0.06) 0.657 (0.08) 0.676 (0.07)
50 0.642 (0.09) 0.583 (0.08) 0.642 (0.09) 0.630 (0.11) 0.676 (0.10) 0.487 (0.16) 0.621 (0.07) 0.674 (0.09) 0.602 (0.10) 0.638 (0.10) 0.583 (0.08) 0.607 (0.10)
31 10 0.696 (0.01) 0.745 (0.04) 0.690 (0.01) 0.534 (0.06) 0.730 (0.03) 0.537 (0.11) 0.710 (0.04) 0.739 (0.04) 0.693 (0.03) 0.680 (0.04) 0.494 (0.02) 0.742 (0.03)
50 0.637 (0.02) 0.679 (0.01) 0.629 (0.03) 0.540 (0.06) 0.657 (0.03) 0.507 (0.02) 0.604 (0.06) 0.643 (0.04) 0.661 (0.03) 0.642 (0.03) 0.531 (0.03) 0.684 (0.02)
37 10 0.510 (0.02) 0.525 (0.03) 0.491 (0.05) 0.524 (0.04) 0.497 (0.01) 0.502 (0.01) 0.554 (0.05) 0.522 (0.02) 0.540 (0.02) 0.530 (0.02) 0.498 (0.03) 0.503 (0.04)
50 0.504 (0.05) 0.489 (0.07) 0.527 (0.05) 0.510 (0.04) 0.502 (0.02) 0.500 (0.01) 0.537 (0.05) 0.488 (0.05) 0.502 (0.04) 0.526 (0.03) 0.494 (0.04) 0.502 (0.04)
23381 10 0.570 (0.06) 0.625 (0.06) 0.570 (0.06) 0.486 (0.02) 0.557 (0.04) 0.538 (0.08) 0.531 (0.16) 0.621 (0.06) 0.555 (0.05) 0.582 (0.05) 0.557 (0.10) 0.615 (0.04)
50 0.581 (0.06) 0.573 (0.05) 0.586 (0.07) 0.506 (0.01) 0.566 (0.13) 0.452 (0.11) 0.435 (0.15) 0.543 (0.02) 0.599 (0.03) 0.499 (0.04) 0.551 (0.06) 0.566 (0.05)
1480 10 0.542 (0.05) 0.534 (0.04) 0.539 (0.04) 0.520 (0.04) 0.547 (0.02) 0.511 (0.05) 0.581 (0.05) 0.534 (0.03) 0.558 (0.03) 0.551 (0.04) 0.518 (0.07) 0.559 (0.06)
50 0.537 (0.07) 0.543 (0.04) 0.531 (0.03) 0.528 (0.04) 0.556 (0.03) 0.479 (0.04) 0.546 (0.05) 0.530 (0.06) 0.519 (0.08) 0.523 (0.08) 0.535 (0.04) 0.546 (0.01)
1067 10 0.802 (0.02) 0.690 (0.04) 0.987 (0.01) 0.957 (0.01) 0.503 (0.01) 0.462 (0.09) 0.486 (0.01) 0.798 (0.11) 0.992 (0.01) 0.993 (0.01) 0.978 (0.01) 0.880 (0.04)
50 0.701 (0.08) 0.600 (0.02) 0.796 (0.03) 0.732 (0.06) 0.510 (0.01) 0.550 (0.09) 0.484 (0.03) 0.568 (0.04) 0.796 (0.04) 0.801 (0.03) 0.792 (0.02) 0.647 (0.09)
334 10 0.643 (0.05) 0.806 (0.10) 0.643 (0.05) 0.767 (0.11) 0.588 (0.05) 0.499 (0.04) 0.622 (0.06) 0.742 (0.17) 0.811 (0.04) 0.767 (0.04) 0.806 (0.10) 0.843 (0.02)
50 0.510 (0.01) 0.575 (0.05) 0.510 (0.01) 0.537 (0.05) 0.504 (0.06) 0.481 (0.03) 0.495 (0.04) 0.510 (0.04) 0.578 (0.02) 0.553 (0.05) 0.575 (0.05) 0.621 (0.07)
1071 10 0.584 (0.08) 0.462 (0.10) 0.526 (0.06) 0.464 (0.10) 0.525 (0.03) 0.483 (0.03) 0.470 (0.04) 0.526 (0.09) 0.536 (0.07) 0.546 (0.07) 0.466 (0.12) 0.430 (0.05)
50 0.612 (0.03) 0.421 (0.11) 0.547 (0.12) 0.442 (0.12) 0.508 (0.01) 0.503 (0.01) 0.448 (0.08) 0.404 (0.06) 0.468 (0.11) 0.445 (0.10) 0.463 (0.07) 0.429 (0.16)
1498 10 0.616 (0.09) 0.633 (0.05) 0.593 (0.09) 0.484 (0.03) 0.632 (0.06) 0.614 (0.10) 0.571 (0.05) 0.617 (0.06) 0.546 (0.05) 0.572 (0.06) 0.512 (0.04) 0.642 (0.06)
50 0.569 (0.06) 0.598 (0.05) 0.556 (0.03) 0.431 (0.06) 0.577 (0.02) 0.457 (0.06) 0.576 (0.05) 0.587 (0.03) 0.536 (0.09) 0.530 (0.08) 0.543 (0.07) 0.602 (0.04)
40691 10 0.555 (0.04) 0.486 (0.03) 0.488 (0.06) 0.497 (0.02) 0.502 (0.03) 0.497 (0.01) 0.500 (0.02) 0.499 (0.03) 0.537 (0.04) 0.539 (0.05) 0.713 (0.01) 0.478 (0.04)
50 0.512 (0.05) 0.506 (0.02) 0.511 (0.01) 0.505 (0.03) 0.500 (0.02) 0.487 (0.01) 0.490 (0.03) 0.499 (0.03) 0.474 (0.06) 0.490 (0.05) 0.642 (0.02) 0.494 (0.03)
181 10 0.524 (0.03) 0.514 (0.01) 0.531 (0.03) 0.521 (0.02) 0.502 (0.04) 0.505 (0.02) 0.510 (0.05) 0.514 (0.01) 0.564 (0.03) 0.571 (0.04) 0.575 (0.01) 0.563 (0.02)
50 0.473 (0.02) 0.507 (0.02) 0.493 (0.02) 0.507 (0.02) 0.502 (0.05) 0.499 (0.01) 0.505 (0.03) 0.496 (0.02) 0.528 (0.03) 0.521 (0.02) 0.522 (0.01) 0.511 (0.02)

### v5

@@ Appendix C Data set-specific AUC classification scores
Dataset ID % Median-GBT Median-FTT MICE-GBT MICE-FTT GAIN-GBT GAIN-FTT Diffputer-GBT Diffputer-FTT MIA-Xgboost MIA-LightGBM AM-FTT IFIAL (K=d/2)
13 10 0.704 (0.08) 0.660 (0.08) 0.704 (0.08) 0.658 (0.06) 0.701 (0.09) 0.698 (0.08) 0.667 (0.08) 0.694 (0.06) 0.550 (0.13) 0.650 (0.06) 0.660 (0.08) 0.675 (0.06)
50 0.642 (0.09) 0.606 (0.08) 0.642 (0.09) 0.638 (0.10) 0.676 (0.10) 0.617 (0.10) 0.628 (0.07) 0.674 (0.09) 0.550 (0.16) 0.633 (0.11) 0.606 (0.08) 0.637 (0.11)
31 10 0.699 (0.01) 0.745 (0.04) 0.695 (0.01) 0.687 (0.04) 0.730 (0.03) 0.697 (0.03) 0.710 (0.04) 0.739 (0.04) 0.582 (0.11) 0.580 (0.06) 0.550 (0.02) 0.708 (0.05)
50 0.641 (0.02) 0.679 (0.01) 0.635 (0.03) 0.644 (0.03) 0.657 (0.03) 0.661 (0.03) 0.618 (0.06) 0.645 (0.04) 0.550 (0.02) 0.573 (0.06) 0.567 (0.03) 0.650 (0.06)
37 10 0.565 (0.02) 0.578 (0.03) 0.550 (0.05) 0.582 (0.02) 0.555 (0.01) 0.590 (0.02) 0.601 (0.05) 0.575 (0.02) 0.559 (0.01) 0.577 (0.04) 0.556 (0.03) 0.805 (0.04)
50 0.559 (0.05) 0.551 (0.07) 0.571 (0.05) 0.570 (0.03) 0.558 (0.02) 0.558 (0.04) 0.576 (0.05) 0.550 (0.05) 0.556 (0.01) 0.562 (0.04) 0.553 (0.04) 0.622 (0.03)
181 10 0.568 (0.03) 0.560 (0.01) 0.573 (0.03) 0.605 (0.04) 0.550 (0.04) 0.600 (0.03) 0.556 (0.05) 0.560 (0.01) 0.552 (0.02) 0.565 (0.02) 0.608 (0.01) 0.742 (0.02)
50 0.550 (0.02) 0.566 (0.02) 0.559 (0.02) 0.572 (0.02) 0.563 (0.05) 0.576 (0.03) 0.565 (0.03) 0.561 (0.02) 0.568 (0.01) 0.566 (0.02) 0.573 (0.01) 0.617 (0.02)
334 10 0.661 (0.05) 0.806 (0.10) 0.661 (0.05) 0.767 (0.04) 0.615 (0.05) 0.811 (0.04) 0.643 (0.06) 0.742 (0.17) 0.550 (0.04) 0.767 (0.11) 0.806 (0.10) 0.744 (0.10)
50 0.558 (0.01) 0.594 (0.05) 0.558 (0.01) 0.582 (0.05) 0.555 (0.06) 0.595 (0.02) 0.550 (0.04) 0.558 (0.04) 0.597 (0.03) 0.573 (0.05) 0.594 (0.05) 0.616 (0.05)
463 10 0.627 (0.09) 0.675 (0.06) 0.596 (0.11) 0.595 (0.10) 0.672 (0.10) 0.665 (0.08) 0.556 (0.08) 0.556 (0.21) 0.627 (0.22) 0.585 (0.12) 0.550 (0.09) 0.557 (0.19)
50 0.653 (0.13) 0.601 (0.12) 0.622 (0.14) 0.554 (0.11) 0.566 (0.09) 0.627 (0.10) 0.663 (0.09) 0.572 (0.11) 0.601 (0.15) 0.550 (0.12) 0.607 (0.21) 0.488 (0.06)
1063 10 0.892 (0.05) 0.671 (0.11) 0.863 (0.04) 0.912 (0.04) 0.572 (0.03) 0.904 (0.04) 0.550 (0.06) 0.595 (0.11) 0.582 (0.02) 0.707 (0.14) 0.797 (0.07) 0.846 (0.06)
50 0.597 (0.06) 0.585 (0.05) 0.564 (0.04) 0.590 (0.08) 0.563 (0.02) 0.588 (0.07) 0.565 (0.08) 0.550 (0.05) 0.568 (0.03) 0.559 (0.05) 0.554 (0.04) 0.705 (0.13)
1067 10 0.802 (0.02) 0.711 (0.04) 0.987 (0.01) 0.993 (0.01) 0.579 (0.01) 0.992 (0.01) 0.567 (0.01) 0.798 (0.11) 0.550 (0.09) 0.957 (0.01) 0.978 (0.01) 0.762 (0.04)
50 0.712 (0.08) 0.636 (0.02) 0.796 (0.03) 0.801 (0.03) 0.569 (0.01) 0.796 (0.04) 0.550 (0.03) 0.613 (0.04) 0.599 (0.09) 0.735 (0.06) 0.792 (0.02) 0.743 (0.02)
1071 10 0.625 (0.08) 0.550 (0.10) 0.589 (0.06) 0.602 (0.07) 0.589 (0.03) 0.595 (0.07) 0.616 (0.04) 0.589 (0.09) 0.624 (0.03) 0.613 (0.10) 0.552 (0.12) 0.690 (0.13)
50 0.612 (0.03) 0.560 (0.11) 0.573 (0.12) 0.572 (0.10) 0.553 (0.01) 0.584 (0.11) 0.574 (0.08) 0.551 (0.06) 0.550 (0.01) 0.571 (0.12) 0.582 (0.07) 0.603 (0.13)
1480 10 0.565 (0.05) 0.561 (0.04) 0.564 (0.04) 0.569 (0.04) 0.568 (0.02) 0.573 (0.03) 0.584 (0.05) 0.561 (0.03) 0.550 (0.05) 0.554 (0.04) 0.553 (0.07) 0.587 (0.02)
50 0.553 (0.07) 0.554 (0.04) 0.552 (0.03) 0.551 (0.08) 0.556 (0.03) 0.550 (0.08) 0.554 (0.05) 0.552 (0.06) 0.552 (0.04) 0.551 (0.04) 0.553 (0.04) 0.556 (0.01)
1498 10 0.632 (0.09) 0.642 (0.05) 0.617 (0.09) 0.604 (0.06) 0.642 (0.06) 0.588 (0.05) 0.604 (0.05) 0.632 (0.06) 0.630 (0.10) 0.550 (0.03) 0.567 (0.04) 0.657 (0.04)
50 0.578 (0.06) 0.598 (0.05) 0.568 (0.03) 0.550 (0.08) 0.583 (0.02) 0.554 (0.09) 0.582 (0.05) 0.590 (0.03) 0.569 (0.06) 0.431 (0.06) 0.559 (0.07) 0.503 (0.07)
23381 10 0.582 (0.06) 0.628 (0.06) 0.582 (0.06) 0.592 (0.05) 0.572 (0.04) 0.570 (0.05) 0.550 (0.16) 0.625 (0.06) 0.556 (0.08) 0.596 (0.02) 0.572 (0.10) 0.642 (0.04)
50 0.595 (0.06) 0.590 (0.05) 0.597 (0.07) 0.550 (0.04) 0.586 (0.13) 0.604 (0.03) 0.570 (0.15) 0.574 (0.02) 0.606 (0.11) 0.554 (0.01) 0.578 (0.06) 0.611 (0.02)
40691 10 0.595 (0.04) 0.550 (0.03) 0.551 (0.06) 0.585 (0.05) 0.560 (0.03) 0.583 (0.04) 0.559 (0.02) 0.560 (0.03) 0.564 (0.01) 0.570 (0.02) 0.713 (0.01) 0.670 (0.03)
50 0.571 (0.05) 0.568 (0.02) 0.571 (0.01) 0.559 (0.05) 0.565 (0.02) 0.550 (0.06) 0.559 (0.03) 0.564 (0.03) 0.557 (0.01) 0.567 (0.03) 0.644 (0.02) 0.647 (0.01)